# Synchronous Consensus in Partial Synchrony

### v1

@@ Abstract.
We demonstrate a deterministic Byzantine consensus algorithm with synchronous performance in partial synchrony and naturally leaderless operation. Each message is authenticated only with the digital signature of its creator and resilience to any number of $f<n/2$ Byzantine processes requires two communication rounds. The algorithm terminates within a bounded interval of time. It is resilient to transient link faults and asynchrony in a fraction of links with a known distinct size per actual number of faulty processes – links asynchrony and faults are circumvented with up to 3-hop epidemic dissemination. Key finding: resilience to asynchrony of links and the enabled by it leaderless consensus ensure algorithm operation with simultaneous validity, safety, and bounded liveness.
@@ 1. Introduction
Consensus in synchrony (Pease et al., 1980) and in partial synchrony (Dwork et al., 1988) are distinctive theoretical domains. The point of separation is a single assumption, based on the finely adjustable per needs definition of communication synchrony (Dwork et al., 1988). Synchronous consensus can tolerate crash fails (Pease et al., 1980), Byzantine faults (Pease et al., 1980) (Dolev and Strong, 1983), and faulty links (Dolev, 1981). Yet once a system accepts that asynchronous operation of some links is occasionally possible, it enters the domain of partial synchrony. Here the processes do not distinguish whether a faulty or an asynchronous link causes the non-receiving of a message. We present a concept and solution for deterministic Byzantine consensus with synchronous performance in partial synchrony. It challenges an established by theory necessity and a conjecture that has never been proved or contested. The established necessity (Dolev and Strong, 1983) states that resilience to $f$ Byzantine processes of a synchronous deterministic consensus protocol using digitally authenticated (Rivest et al., 1978) messages requires $(f+1)$ communication rounds. The conjecture (Gilbert and Lynch, 2012) states that no consensus protocol can operate in conditions of partial synchrony with ensured simultaneous safety, validity, and liveness. When applied to leader-based protocols, the conjecture looks axiomatic. Yet the case is different with bounded delivery in spite of asynchrony in a fraction of links. It naturally makes the long aspired, but never solved for practical purposes, leaderless consensus (Boran and Schiper, 2010) (Antoniadis et al., 2021) in partial synchrony a no more existing problem.
@@ 1.1. Background
The term partial synchrony reflects the practice-driven attention to communication for consensus that is neither fully asynchronous nor fully synchronous. As a theoretical domain between synchrony and asynchrony, it can be viewed as two separate theoretical spaces with a borderline blurred by the conditionality of tolerated asynchrony.
Non-manageable space: Until recently the only approach in dealing with asynchrony was the model of eventual synchrony (Dwork et al., 1988). It reflects the reality where intervals of non-infringed synchrony with unknown duration are followed by intervals with infringed synchrony. It accepts the inevitability of infringed synchrony and allows termination only during large enough intervals with non-infringed synchrony (Dutta and Guerraoui, 2002) (Alistarh et al., 2008). Yet, a prepublication (Klianev, 2023) shows solving consensus in complete asynchrony under the model used for proving the FLP impossibility result (Fischer et al., 1985), on condition of non-dependence of agreement on the content of the initial values.
Manageable space: The preferred by industry model (Castro and Liskov, 2002) is to rely on a consensus leader rather than wait for eventual synchrony. With this model, the dependence on asynchrony is replaced with dependence on a possibly Byzantine (Lamport et al., 1982) single point of failure. The use of cryptographically authenticated messages (Rivest et al., 1978) prevents Byzantine processes from doing more harm on consensus safety than the crash-failed processes and allows solving consensus with a simple majority of processes (Pease et al., 1980) (Dolev and Strong, 1983). Nevertheless it cannot prevent the ability of a Byzantine leader to spoil liveness.
None of the above models ensures timely termination in the presence of a manageable level of asynchrony. With the model of eventual synchrony, the eventuality is contingent in time as it depends on external factors. With a model relying on a leader, a Byzantine leader can purposefully delay the broadcasting of its consensus-value proposing message without being detected. We apply gossiping (Birman et al., 1999) (Eugster et al., 2004), normally used to circumvent faulty and non-existing links, for bounded delivery in presence of asynchrony and demonstrate that leaderless consensus and bounded termination are natural outcomes from the ability to mitigate some amount of communication asynchrony.
@@ 1.2. Motivation
In contrast to vulnerable (Bernstein et al., 1987) distributed transactions (Lindsay et al., 1979) (Gray, 1978) (Lampson and Sturgis, 1976) across centralised databases, the significance of decentralised forms of distributed data systems steadily grows and shows no signs of fading. These data systems minimise the necessary level of trust between participants and ability of one participant to exercise control over another. They make reconciliation between transacting parties with data stored in a single decentralised silo much easier, and reduce their own points of weakness by eliminating reliance on specific actors. These systems operate with global replication of local transactions and ensure atomic (linearisable) (Herlihy and Wing, 1990) consistency across local systems with state machine replication (Lamport, 1978) (Schneider, 1990) (Lampson, 1996), based on sequentially ordered requests for local transactions. Agreement for the purposes of state machine replication is about an ordered set of initial data and its actual values cannot make termination impossible (Klianev, 2023). Decentralised systems, performing with time-bounded liveness and no single point of failure, can interoperate with transactions coordinated by participants with a time-controlled workflow of steps (Garcia-Molina and Salem, 1987).
@@ 1.3. Problem Formulation
We demonstrate solving a problem with the following formulation: Given: i) consensus system $S$ of $n$ processes, $n>2f$, where $f$ is the tolerated number of crash-failed and Byzantine processes; ii) communicate using authenticated messages; iii) every process $P_{i}$, where $i\in\{1,...,n\}$, has an initial value $v_{i}$; iv) set $C\subseteq S$ comprises no less than $(n-f)$ correct processes; v) every $P_{i}\in C$ delivers own message $m_{i}$ to every other $P\in C$ over synchronous links with up to 3 hops during periods with asynchronous and/or faulty links; and vi) terminates with set of processes $C^{\prime}$, $(C^{\prime}\subseteq C,|C^{\prime}|\geq(n-f))$. Demonstrate: ability to ensure that all processes in $C^{\prime}$ agree on an $n$-vector $V=(v_{1},...,v_{n})$ that contains the initial values of at least all processes in $C^{\prime}$, and less than $f$ vector elements are replaced with $\varnothing$ denoting an empty element, with 2 rounds of exchange of messages and within a time interval $T_{T}\leq 2*3*T_{S}$, where $T_{S}$ is the time boundary for synchrony of links.
@@ 1.4. Contribution
1. Quantified tolerance to links asynchrony and faults. 2. $f$-resilience, $f<n/2$, implemented with 2 messaging rounds. 3. Ensured simultaneous safety, validity, and bounded liveness. 4. Naturally leaderless operation.
@@ 1.5. Similar Works
Synchronous Consensus with Transient Faults Santoro et al (Santoro and Widmayer, 1989) studies the possibility for agreement in a fully synchronous system and presents conditions making the agreement impossible. It shows that $(n/2+1)$ binary agreement is impossible if, at each time unit, there is one process whose all $(n-1)$ messages are lost. This result applies to consensus protocols aiming agreement that depends on the content of the initial values. It does not apply to protocols where the agreement does not depend on the content of the initial values, such as agreement about ordered initial values. Synchronous Consensus Over a Fraction of Links Siu et al (Siu et al., 1998) presents a concept for fault-tolerant virtual channels in a synchronous deterministic Byzantine consensus system, where processes and links can fail independently and simultaneously. It ensures reliable delivery for the purposes of synchronous consensus over a fraction of links but does not consider links asynchrony. Aguilera et al (Aguilera et al., 2003) (Aguilera et al., 2004) demonstrate a case of solved synchronous deterministic Byzantine consensus with a small fraction of all links operating synchronously. The shown result – one correct process having all inbound and outbound links synchronous – is with a negligibly small probability to happen. It shows an exception from the normal distribution of asynchrony rather than a solution.
Consensus via Gossip-based Reliable Multicast In a distributed system where every process knows every other, the algorithm of Birman et al (Birman et al., 1999) uses gossips, a.k.a. epidemic dissemination of messages, to deliver the messages to all recipients, even if the sender fails before sending to all. The algorithm tolerates process crash failures and a bounded number of failures of links. Yet this algorithm does not consider Byzantine failures, asynchrony of links, or bounded delivery. Atomic Broadcast for Synchronous Replicated Storage Cristian et al (Cristian et al., 1995) demonstrates ability to handle random communication delays and link failures, using multi-hop delivery in presence of crashed and Byzantine processes, with objective of information diffusion for the purposes of synchronous replicated storage. Even so, it does not show ensured bounded termination. In contrast, we use multi-hop delivery to: i) mitigate the effect of asynchronous and/or faulty links; ii) neutralize with reliable multicast the effect of current-round crashes and Byzantine faults; and iii) ensure time-bounded termination of consensus algorithm.
@@ 2.1. Assumptions
Tolerance Limits: A plain majority of correct processes with bounded indirect delivery of messages between themselves must exist at any time with no reliance on asynchronous links becoming synchronous, or faulty links becoming correct, within the current consensus round.
@@ 2.2. The Model
$R^{Out}=(R^{In}_{1},R^{In}_{2},...,R^{In}_{n})$ Internal state of a process $P$ comprises value in $R^{In}$ and value in $R^{Out}$, together with a program counter and internal storage.
Initial values of the internal state are fixed, with an exception for the initial value in $R^{In}$. More specifically, $R^{In}$ starts with a length prefixed string having an arbitrary value from a finite universe, and $R^{Out}$ starts always with the fixed value $R^{Out}=(\varnothing,\varnothing,...,\varnothing)$, where $\varnothing$ is a null marker denoting an empty vector element. Objective of the consensus protocol is to establish agreement about the value of $R^{Out}=(R^{In}_{1},R^{In}_{2},...,R^{In}_{n})$ between $\geq(n-f)$ correct processes with exchange of two type of messages: - Initial Value message of $P_{i}$ contains the initial value in $R^{In}_{i}$;
@@
A message is a tuple $(P_{s},P_{d},m)$, where $P_{s}$ is the ’source’ process, $P_{d}$ is the ’destination process’, and $m$ is a ’message value’ from a fixed universe $M$. The message system maintains a message buffer for each one-way link $L_{pq}$ that delivers messages from process $P_{p}$ to process $P_{q}$. Message buffer is a multi-set of messages that have been sent but not yet delivered, in case it models the behaviour of a correct synchronous or a correct asynchronous link. It is an empty set of messages in case it models a faulty link. The message system supports two abstract operations: - Send($P_{s},P_{d},m$). It places ($P_{s},P_{d},m$) in a message buffer. - Receive($P_{d}$). It deletes some message ($P_{s},P_{d},m$) from the buffer and returns $m$, in which case ($P_{s},P_{d},m$) is delivered, or $\varnothing$ and leaves the buffer unchanged. A message buffer representing a link $L_{pq}$ operates with a delay made by two components: one representing the time necessary for a message to physically travel from process $P_{p}$ to process $P_{q}$; and another representing the delay resulting from congestions on the physical link. The delay operates per message by loading a counter with the delay value on every message entering the buffer. Receive($P_{q}$) does not return $\varnothing$ only if among all buffers representing links to process $P_{q}$ there is a message with expired delay. Thus, the message system acts deterministically. Every message ($P_{p},P_{q},m$) in the message buffer is delivered, subject only to the condition that Receive($P_{d}$) is performed until expiration of delay $D_{pq}$ associated with link $L_{pq}$ at the moment the message had entered the buffer. Link $L_{pq}$ is considered synchronous if $D_{pq}$ is not larger than the delivery time considered synchronous and asynchronous otherwise. Message Event is a tuple ($P_{s},P_{d},m$) indicating that $m$, having content created by process $P_{s}$, has been delivered to process $P_{d}$. Time Event is a trigger for execution of a sequence of activities related to preparation and broadcasting of a protocol message. Decision value is a vector $V^{De}$ of $n$ elements that contains no more than $f$ elements with value $\varnothing$. Decision state is an internal state a process $P$ reaches after $P$ knows that no less than $(n-f)$ processes, including itself, will propose exactly the same $A^{Prop}$ for decision value $V^{De}$. Once $P$ has reached a decision state, $A^{Prop}$ becomes $V^{De}$ and $P$ writes $V^{De}$ in its output register $R^{Out}$.
@@
- Correct process. A process is correct in a run if it takes infinitely many steps and strictly follows function $\Phi$. - Crash-failed process. A process that executes correctly until unexpectedly stops and once stopped, cannot restart. - Byzantine-faulty process. A process that deliberately deviates from the strict following of function $\Phi$.
@@
- Decision. A decision configuration is an accessible configuration where at least $(n-f)$ processes have reached a valid decision state. - Valid decision state. A vector with size $n$ that contains the initial values of no less than the agreed processes as vector elements and $\varnothing$ marker as value in the rest of vector elements. Run types:
@@ 3. The Solution
The consensus algorithm we are going to present consists of two parts: i) indirect delivery; and ii) synchronous Byzantine consensus protocol with agreement about ordered initial values (Pease et al., 1980).

### v3

@@ Abstract.
We demonstrate a deterministic Byzantine consensus algorithm with synchronous operation in partial synchrony. It is naturally leaderless, tolerates any number of $f<n/2$ Byzantine processes with 2 rounds of exchange of originator-only signed messages, and terminates within a bounded interval of time. The algorithm is resilient to transient faults and asynchrony in a fraction of links with known size per number of faulty processes. It circumvents asynchronous and faulty links with 3-hop epidemic dissemination. Key finding: the resilience to asynchrony of links and the enabled by it leaderless consensus in partial synchrony ensure algorithm operation with simultaneous validity, safety, and bounded liveness.
@@ 1. Introduction
The theoretical divergence between consensus in synchrony (Pease et al., 1980) and in partial synchrony (Dwork et al., 1988) is to a certain degree not natural. Point of separation is a single assumption, based on the definition of communication synchrony (Dwork et al., 1988), which allows tuning per needs. Synchronous consensus can tolerate crash fails (Pease et al., 1980), Byzantine faults (Pease et al., 1980) (Dolev and Strong, 1983), and faulty links (Dolev, 1981), but not asynchronous links. This is despite the observations that asynchronous links have the same effect on synchronous consensus as the faulty links.
We solved the impossibility of bounded consensus termination in partial synchrony. The number of tolerated asynchronous links is a function of the number of faulty processes per system size. The asynchronous links are circumvented with 3-hop epidemic dissemination (Birman et al., 1999) and thus the time for delivery is 3 times longer, yet bounded and known in advance. Critical for the solution is its consensus algorithm where the number of communication rounds, needed for resilience of synchronous deterministic protocol with authenticated messages (Rivest et al., 1978) to f Byzantine processes, is smaller than the known minimum of $(f+1)$ communication rounds (Dolev and Strong, 1983). The algorithm uses only 2 communication rounds regardless of the number of tolerated Byzantine processes. With more than 12 processes in partial synchrony, it terminates faster than a synchronous Byzantine (Dolev and Strong, 1983) system with the same size and delivery bound. Moreover, it makes leaderless consensus (Boran and Schiper, 2010) (Antoniadis et al., 2021) in partial synchrony a no more existing problem.
@@ 1.1. Background
The term partial synchrony reflects the practice-driven attention to messaging that is neither fully asynchronous nor fully synchronous. As a theoretical field, it can be viewed as two separate spaces with a borderline blurred by the conditionality of tolerated asynchrony.
Unmanageable space: Until recently the only approach in dealing with asynchrony was the model of eventual synchrony (Dwork et al., 1988). It reflects the reality where intervals of non-infringed synchrony with unknown duration are followed by intervals with infringed synchrony. It accepts the inevitability of infringed synchrony and allows termination only during large enough intervals with non-infringed synchrony (Dutta and Guerraoui, 2002) (Alistarh et al., 2008). Yet, a prepublication (Klianev, 2023) shows solving consensus in complete asynchrony under the model used for proving the FLP impossibility result (Fischer et al., 1985), on condition of non-dependence of agreement on the content of the initial values.
Manageable space: The industry-preferred model (Castro and Liskov, 2002) is to rely on a consensus leader rather than wait for eventual synchrony. With this model, the dependence on asynchrony is replaced with dependence on a possibly Byzantine (Lamport et al., 1982) single point of failure. The use of cryptographically authenticated messages (Rivest et al., 1978) prevents Byzantine processes from doing more harm on consensus safety than the crash-failed processes and allows solving consensus with a simple majority of processes (Pease et al., 1980) (Dolev and Strong, 1983). Nevertheless it cannot prevent the ability of a Byzantine leader to spoil liveness.
None of the above models ensures bounded termination in the presence of asynchrony. The termination either depends on external factors beyond control or on a possibly Byzantine leader that can delay the broadcasting of its proposing message without being detected. We apply gossiping (Birman et al., 1999) (Eugster et al., 2004), normally used to circumvent faulty and non-existing links, for bounded delivery in presence of asynchrony and show that leaderless consensus and bounded termination are natural outcomes from the ability to mitigate some amount of asynchrony.
@@ 1.2. Motivation
The decentralised forms of distributed data systems operate with global replication of local transactions and atomic (Herlihy and Wing, 1990) consistency across local systems via state machine replication (Lamport, 1978) (Schneider, 1990) (Lampson, 1996). Agreement for the purposes of state machine replication is about an ordered set of initial data and its actual values cannot make consensus impossible (Klianev, 2023). Systems with no single point of failure and bounded liveness could solve the vulnerability (Bernstein et al., 1987) of distributed transactions (Lindsay et al., 1979) (Gray, 1978) (Lampson and Sturgis, 1976) via being coordinated directly by its participants with a time-controlled workflow (Garcia-Molina and Salem, 1987).
Moreover, the possibility for bounded termination of consensus in partial synchrony is expected to trigger attention on the so far axiomatic acceptance that a consensus protocol cannot operate with ensured simultaneous validity, safety, and liveness in partial synchrony (Gilbert and Lynch, 2012), and even less with a bounded liveness.
@@ 1.3. Problem Formulation
We have to show solving a problem with the following formulation:
Given: i) consensus system $S$ of $n$ processes, $n>2f$, where $f$ is the tolerated number of crash-failed and Byzantine processes; ii) communicate using authenticated messages; iii) every process $P_{i}$, where $i\in\{1,...,n\}$, has an initial value $v_{i}$; iv) set $C\subseteq S$ comprises no less than $(n-f)$ correct processes; v) every $P_{i}\in C$ delivers own message $m_{i}$ to every other $P\in C$ over synchronous links with up to 3 hops during periods with asynchronous and/or faulty links; and vi) terminates with set of processes $C^{\prime}$, $(C^{\prime}\subseteq C,|C^{\prime}|\geq(n-f))$.
Demonstrate: ability to ensure that all processes in $C^{\prime}$ agree on an $n$-vector $V=(v_{1},...,v_{n})$ that contains the initial values of at least all processes in $C^{\prime}$, and less than $f$ vector elements are replaced with $\varnothing$ denoting an empty element, with 2 rounds of exchange of messages and within a time interval $T_{T}\leq 2*3*T_{S}$, where $T_{S}$ is the time bound for synchrony of links.
@@ 1.4. Contribution
1. Quantified tolerance to links asynchrony and faults.
2. $f$-resilience, $f<n/2$, implemented with 2 messaging rounds.
3. Ensured simultaneous validity, safety, and bounded liveness.
4. Naturally leaderless operation.
@@ 1.5. Similar Works
Synchronous Consensus with Transient Faults
Santoro et al (Santoro and Widmayer, 1989) studies the possibility for agreement in a fully synchronous system and presents conditions making the agreement impossible. It shows that $(n/2+1)$ binary agreement is impossible if, at each time unit, there is one process whose all $(n-1)$ messages are lost. This result applies to agreements that depend on the content of the initial values. It does not apply (Klianev, 2023) to our agreement, which is about ordered initial values.
Synchronous Consensus Over a Fraction of Links
Siu et al (Siu et al., 1998) presents a concept for fault-tolerant virtual channels in a synchronous deterministic Byzantine consensus system, where processes and links can fail independently and simultaneously. It ensures reliable delivery for the purposes of synchronous consensus over a fraction of links but does not consider links asynchrony.
Aguilera et al (Aguilera et al., 2003) (Aguilera et al., 2004) demonstrate a case of solved synchronous deterministic Byzantine consensus with a small fraction of all links operating synchronously. The shown result – one correct process having all inbound and outbound links synchronous – is with a negligibly small probability to happen and it is not a solution.
Consensus via Gossip-based Reliable Multicast
In a distributed system where every process knows every other, the algorithm of Birman et al (Birman et al., 1999) uses gossips, a.k.a. epidemic dissemination of messages, to deliver the messages to all recipients, even if the sender fails before sending to all. The algorithm tolerates process crash failures and a number of failures of links. Yet it not about Byzantine failures, asynchrony of links, or bounded delivery.
Atomic Broadcast for Synchronous Replicated Storage
Cristian et al (Cristian et al., 1995) demonstrates ability to handle random communication delays and link failures, using multi-hop delivery in presence of crashed-failed and Byzantine processes, with objective of information diffusion for the purposes of synchronous replicated storage. Even so, it does not show ensured bounded termination. In contrast, we use multi-hop delivery to mitigate the effect of asynchronous and/or faulty links and ensure bounded termination.
@@ 2.1. Assumptions
Tolerance Limits: A plain majority of correct processes with bounded indirect delivery between themselves exists at any time with no reliance on asynchronous links becoming synchronous, or faulty links becoming correct, within a consensus round.
@@ 2.2. The Model
$R^{Out}=(R^{In}_{1},R^{In}_{2},...,R^{In}_{n})$
Internal state of a process $P$ comprises value in $R^{In}$ and value in $R^{Out}$, together with a program counter and internal storage.
Initial values of the internal state are fixed, with an exception for the initial value in $R^{In}$. More specifically, $R^{In}$ starts with a length prefixed string having an arbitrary value from a finite universe, and $R^{Out}$ starts always with the fixed value $R^{Out}=(\varnothing,\varnothing,...,\varnothing)$, where $\varnothing$ is a null marker denoting an empty vector element.
Objective of the consensus protocol is to establish agreement about the value of $R^{Out}=(R^{In}_{1},R^{In}_{2},...,R^{In}_{n})$ between $\geq(n-f)$ correct processes with exchange of two type of messages:
- Initial Value message of $P_{i}$ contains the initial value in $R^{In}_{i}$;
@@
A message is a tuple $(P_{s},P_{d},m)$, where $P_{s}$ is the ’source’ process, $P_{d}$ is the ’destination process’, and $m$ is a ’message value’ from a fixed universe $M$. The message system maintains a message buffer for each one-way link $L_{pq}$ that delivers messages from process $P_{p}$ to process $P_{q}$. Message buffer is a multi-set of messages that have been sent but not yet delivered, in case it models the behaviour of a correct synchronous or a correct asynchronous link. It is an empty set of messages in case it models a faulty link. The message system supports two abstract operations:
- Send($P_{s},P_{d},m$). It places ($P_{s},P_{d},m$) in a message buffer.
- Receive($P_{d}$). It deletes some message ($P_{s},P_{d},m$) from the buffer and returns $m$, in which case ($P_{s},P_{d},m$) is delivered, or $\varnothing$ and leaves the buffer unchanged.
A message buffer representing a link $L_{pq}$ operates with a delay made by two components: one representing the time necessary for a message to physically travel from process $P_{p}$ to process $P_{q}$; and another representing the delay resulting from congestions on the physical link. The delay operates per message by loading a counter with the delay value on every message entering the buffer. Receive($P_{q}$) does not return $\varnothing$ only if among all buffers representing links to process $P_{q}$ there is a message with expired delay. Thus, the message system acts deterministically. Every message ($P_{p},P_{q},m$) in the message buffer is delivered, subject only to the condition that Receive($P_{d}$) is performed until expiration of delay $D_{pq}$ associated with link $L_{pq}$ at the moment the message had entered the buffer. Link $L_{pq}$ is considered synchronous if $D_{pq}$ is not larger than the delivery time considered synchronous and asynchronous otherwise.
Message Event is a tuple ($P_{s},P_{d},m$) indicating that $m$, having content created by process $P_{s}$, has been delivered to process $P_{d}$.
Time Event is a trigger for execution of a sequence of activities related to preparation and broadcasting of a protocol message.
Decision value is a vector $V^{De}$ of $n$ elements that contains no more than $f$ elements with value $\varnothing$. Decision state is an internal state a process $P$ reaches after $P$ knows that no less than $(n-f)$ processes, including itself, will propose exactly the same $A^{Prop}$ for decision value $V^{De}$. Once $P$ has reached a decision state, $A^{Prop}$ becomes $V^{De}$ and $P$ writes $V^{De}$ in its output register $R^{Out}$.
@@
- Correct process. A process is correct in a run if it takes infinitely many steps and strictly follows function $\Phi$.
- Crash-failed process. A process that executes correctly until unexpectedly stops and once stopped, cannot restart.
- Byzantine-faulty process. A process that deliberately deviates from the strict following of function $\Phi$.
@@
- Decision. A decision configuration is an accessible configuration where at least $(n-f)$ processes have reached a valid decision state.
- Valid decision state. A vector with size $n$ that contains the initial values of no less than the agreed processes as vector elements and $\varnothing$ marker as value in the rest of vector elements.
Run types:
@@ 3. The Solution
The consensus algorithm consists of two parts: i) indirect delivery; and ii) synchronous Byzantine consensus protocol with agreement about ordered initial values (Pease et al., 1980).